# The Stepanov theorem for Q-valued functions

### v1

@@ 3. Q-valued functions
On a generic metric space $(M,d)$ we can define the space of $Q$-points $\mathcal{A}_{Q}\left(M\right)$, where $Q$ is a positive integer, as the set of all measures $\mathbf{P}$ on $M$ that can be represented as the sum of $Q$ Dirac measures $\left\llbracket P_{1}\right\rrbracket+\left\llbracket P_{2}\right\rrbracket+\dotsb+\left\llbracket P_{Q}\right\rrbracket$ not necessarily distinct. Then a $Q$-valued function from a set $A$ to $M$ is just a function defined on $A$ with values on $\mathcal{A}_{Q}\left(M\right)$, clearly $1$-valued functions from $A$ to $M$ are equivalent to classical (single-valued) functions from $A$ to $M$.
Moreover, $Q$ single valued functions $f_{1},f_{2},\dotsc,f_{Q}$ can be used to compose a single $Q$-valued function $f(x)=\left\llbracket f_{1}(x)\right\rrbracket+\left\llbracket f_{2}(x)\right\rrbracket+\dotsb+\left\llbracket f_{Q}(x)\right\rrbracket$, converselt every $Q$-valued function $f$ can be globally decomposed into $Q$ single valued functions by choosing a total ordering on $M$. However, such decomposition is never useful except when $M$ is a really good space (usually $M={\mathbb{R}}$) because many useful properties of $f$ will not be inherited by such general decompositions.
@@
$\mathcal{G}\left[\mathbf{A},\mathbf{B}\right]=\mathcal{G}\left[\sum^{Q}_{i=1}\left\llbracket A_{i}\right\rrbracket,\sum^{Q}_{i=1}\left\llbracket B_{i}\right\rrbracket\right]=\min_{\sigma\in\mathcal{S}_{Q}}\sqrt{\sum^{Q}_{i=1}d\left(A_{1},B_{\sigma(i)}\right)^{2}}$
where $\mathcal{S}_{Q}$ is the set of all the permutations on $\left\{{1,2,\dotsc,Q}\right\}$. Therefore a $Q$-valued function $f$ from a topological/metric/measurable space $\Omega$ to $M$ is continuous/Lipschitz/measurable if and only if ${f}:{\Omega}\rightarrow{\mathcal{A}_{Q}\left(M\right)}$ is continuous/Lipschitz/measurable. Every measurable $Q$-valued function can be globally decomposed into $Q$ measurable single-valued functions (see [3]), however continuous $Q$-valued functions cannot usually be decomposed into $Q$ continuous single-valued functions, for example the following function
@@
Notice that $\mathcal{A}_{Q}\left(M\right)$ is always a complete metric space whenever $M$ is complete, and can be proved by induction on $Q$. This is clearly trivial for $Q=1$, thus assume both $M$ and $\mathcal{A}_{Q}\left(M\right)$ complete. Let $\mathbf{P}_{l}\in\mathcal{A}_{Q+1}\left(M\right)$ be a Cauchy sequence, by induction on $l$ we can easily find a subsequence in the form $\mathbf{P}_{h_{l}}=\left\llbracket P_{l}\right\rrbracket+\mathbf{P}^{\prime}_{l}$, where both $P_{l}\in M$ and $\mathbf{P}^{\prime}_{l}\in\mathcal{A}_{Q}\left(M\right)$ are Cauchy sequences that converge to some $P\in M$ and $\mathbf{P}^{\prime}\in\mathcal{A}_{Q}\left(M\right)$ respectively. Therefore $\mathbf{P}_{h_{l}}\to\left\llbracket P\right\rrbracket+\mathbf{P}^{\prime}$ which in turn implies $\mathbf{P}_{l}\to\left\llbracket P\right\rrbracket+\mathbf{P}^{\prime}$ because $\mathbf{P}_{l}$ is Cauchy.
Many authors (see for example [3]) imposes condition (3.2) in their definition of differentiable $Q$-valued functions since it makes easier to work componentwise. For example if $M={\mathbb{R}}^{n}$ and $\left[\mathbf{F}\right]\in\mathcal{D}^{Q}$ is a differential germ for $f=\sum^{Q}_{i=1}\left\llbracket f_{i}\right\rrbracket$ at some $x\in{\mathbb{R}}^{k}$ with $f_{i}$ continuous at $x$ then we can always take the $i$-th component $F_{i}\in\mathcal{D}$ of $\mathbf{F}$ so that $\left[F_{i}\right]$ is a $\mathcal{D}$-differential for $f_{i}$ at $x$.
Instead other authors (for example [7]) do not use (3.2) in their definition of differentiable $Q$-valued functions, since with (3.2) we could have a sequence of $C^{1}$ $Q$-valued functions $f_{n}$ such that both $f_{n}$ and $\nabla f_{n}$ uniformly converge to some functions $f$ and $F$ respectively but $f$ is not differentiable everywhere! Therefore when one need to work with sequences of $C^{1}$ $Q$-valued functions it may be useful to require only (3.1) without (3.2).
However, if one need only a.e. differentiability (for example when dealing with Sobolev functions) then (3.2) is unimportant as we will see later. To show this, in this place we will always assume that germs in $\mathcal{D}^{Q}$ satisfy both (3.1) and (3.2).
@@
where $\eta\in S_{Q}$ with $\eta(i)=i$ for every $1\leq i\leq R$. We will prove only the last assertion. First of all assume that
@@
for every $y\in\Lambda$.
Combining theorem 3.3 with theorem 2.12 we finally get the following generalized version of Stepanov’s theorem for $Q$-valued functions which generalizes theorem 1.2:
Theorem 3.4.
Let $(X,\mathtt{J},\mu)$ be a doubling metric measure space, $(M,d)$ be a complete metric space and $\mathcal{D}$ a complete (approximate) $X$-distribution on $M$. Then every $Q$-valued function ${f}:{X}\rightarrow{\mathcal{A}_{Q}\left(M\right)}$ is a.e. (approximately) $\mathcal{D}^{Q}$-differentiable on $A_{f}$ (or $A_{f}^{\mu}$).
In particular, these properties hold whenever $M={\mathbb{R}}^{k}$ and $X$ is ${\mathbb{R}}^{n}$ or a $p$-Poincaré doubling metric measure space thanks to [2].

### v3

@@ 3. Q-valued functions
On a generic metric space $(M,d)$ we can define the space of $Q$-points $\mathcal{A}_{Q}\left(M\right)$, where $Q$ is a positive integer, as the set of all measures $\mathbf{P}$ on $M$ that can be represented as the sum of $Q$ Dirac measures $\left\llbracket P_{1}\right\rrbracket+\left\llbracket P_{2}\right\rrbracket+\dotsb+\left\llbracket P_{Q}\right\rrbracket$ not necessarily distinct. Then a $Q$-valued function from a set $A$ to $M$ is just a function defined on $A$ with values on $\mathcal{A}_{Q}\left(M\right)$. Clearly, $1$-valued functions from $A$ to $M$ are equivalent to classical (single-valued) functions from $A$ to $M$.
Moreover, $Q$ single valued functions $f_{1},f_{2},\dotsc,f_{Q}$ can be used to compose a single $Q$-valued function $f(x)=\left\llbracket f_{1}(x)\right\rrbracket+\left\llbracket f_{2}(x)\right\rrbracket+\dotsb+\left\llbracket f_{Q}(x)\right\rrbracket$, conversely every $Q$-valued function $f$ can be globally decomposed into $Q$ single valued functions by choosing a total ordering on $M$. However, such decomposition is never useful except when $M$ is a really good space (usually $M={\mathbb{R}}$) because many useful properties of $f$ will not be inherited by such general decompositions.
@@
$\mathcal{G}\left[\mathbf{A},\mathbf{B}\right]=\mathcal{G}\left[\sum^{Q}_{i=1}\left\llbracket A_{i}\right\rrbracket,\sum^{Q}_{i=1}\left\llbracket B_{i}\right\rrbracket\right]=\min_{\sigma\in\mathcal{S}_{Q}}\sqrt{\sum^{Q}_{i=1}d\left(A_{i},B_{\sigma(i)}\right)^{2}}$
where $\mathcal{S}_{Q}$ is the set of all the permutations on $\left\{{1,2,\dotsc,Q}\right\}$. Therefore a $Q$-valued function $f$ from a topological/metric/measurable space $\Omega$ to $M$ is continuous/Lipschitz/measurable if and only if ${f}:{\Omega}\rightarrow{\mathcal{A}_{Q}\left(M\right)}$ is continuous/Lipschitz/measurable. Every measurable $Q$-valued function can be globally decomposed into $Q$ measurable single-valued functions (see [DeLellis2011]), however continuous $Q$-valued functions cannot usually be decomposed into $Q$ continuous single-valued functions, for example the following function
@@
Notice that $\mathcal{A}_{Q}\left(M\right)$ is always a complete metric space whenever $M$ is complete, and can be proved by induction on $Q$. This is trivial for $Q=1$, thus assume both $M$ and $\mathcal{A}_{Q}\left(M\right)$ are complete. Let $\mathbf{P}_{l}\in\mathcal{A}_{Q+1}\left(M\right)$ be a Cauchy sequence, by induction on $l$ we can easily find a subsequence in the form $\mathbf{P}_{h_{l}}=\left\llbracket P_{l}\right\rrbracket+\mathbf{P}^{\prime}_{l}$, where both $P_{l}\in M$ and $\mathbf{P}^{\prime}_{l}\in\mathcal{A}_{Q}\left(M\right)$ are Cauchy sequences that converge to some $P\in M$ and $\mathbf{P}^{\prime}\in\mathcal{A}_{Q}\left(M\right)$ respectively. Therefore $\mathbf{P}_{h_{l}}\to\left\llbracket P\right\rrbracket+\mathbf{P}^{\prime}$ which in turn implies $\mathbf{P}_{l}\to\left\llbracket P\right\rrbracket+\mathbf{P}^{\prime}$ because $\mathbf{P}_{l}$ is Cauchy.
Many authors (see for example [DeLellis2011]) impose condition (3.2) in their definition of differentiable $Q$-valued functions since it makes it easier to work componentwise. For example if $M={\mathbb{R}}^{n}$ and $\left[\mathbf{F}\right]\in\mathcal{D}^{Q}$ is a differential germ for $f=\sum^{Q}_{i=1}\left\llbracket f_{i}\right\rrbracket$ at some $x\in{\mathbb{R}}^{k}$ with $f_{i}$ continuous at $x$ then we can always take the $i$-th component $F_{i}\in\mathcal{D}$ of $\mathbf{F}$ so that $\left[F_{i}\right]$ is a $\mathcal{D}$-differential for $f_{i}$ at $x$.
Instead other authors (for example [Simon2016]) do not use (3.2) in their definition of differentiable $Q$-valued functions, since with (3.2) we could have a sequence of $C^{1}$ $Q$-valued functions $f_{n}$ such that both $f_{n}$ and $\nabla f_{n}$ uniformly converge to some functions $f$ and $F$ respectively but $f$ is not differentiable everywhere. Therefore when one needs to work with sequences of $C^{1}$ $Q$-valued functions it may be useful to require only (3.1) without (3.2).
However, if one needs only a.e. differentiability for functions from ${\mathbb{R}}^{n}$ into $\mathcal{A}_{Q}\left({\mathbb{R}}^{k}\right)$ with respect to the Lebesgue measure on ${\mathbb{R}}^{n}$ (for example when dealing with Sobolev functions) then (3.2) is unimportant because the set of differentiability points for which (3.2) is not satisfied has zero measure. In this place we will always assume that germs in $\mathcal{D}^{Q}$ satisfy both (3.1) and (3.2).
@@
where $\eta\in S_{Q}$ with $\eta(i)=i$ for every $1\leq i\leq R$. We will prove only the last assertion. First of all, assume that
@@
for every $y\in\Lambda$. ∎
With this result we are finally able to prove theorem 1.3.
Proof of theorem 1.3.
Let $\mathcal{D}$ be a complete $X$-distribution for $(M,d)$, then by theorem 3.3 the $X$-distribution $\mathcal{D}^{Q}$ for $\left(\mathcal{A}_{Q}\left(M\right),\mathcal{G}\right)$ is complete for every $Q\geq 1$. Then by theorem 2.12 we immediately get that every function ${f}:{X}\rightarrow{\mathcal{A}_{Q}\left(M\right)}$ is a.e. $\mathcal{D}^{Q}$-differentiable on $A_{f}$. ∎
[…]